# RichRAG: Crafting Rich Responses for Multi-faceted Queries in Retrieval-Augmented Generation

### v2

@@ 1 Introduction
Though some advanced techniques Jiang et al. (2023); Asai et al. (2024); Wang et al. (2023c); Li et al. (2024) have been proposed, existing studies primarily focus on addressing specific problems that require concise and definitive answers. However, user intents are complex and multi-faceted, necessitating rich and comprehensive answers. As Figure 1 shows, when a user inquires about rapping-related information, a rich response about various aspects of rapping, such as origins, characteristics, and evolution could lead to a more satisfactory user experience than a superficial description.
@@
Our research is focused on developing effective RAG approaches to handle these more complex user needs. We propose a RAG framework, RichRAG, which is designed to offer diverse external knowledge that comprehensively covers various sub-aspects of multi-faceted queries, thereby enhancing the downstream generator (an LLM) to yield rich responses. RichRAG first employs a sub-aspect explorer to explicitly predict sub-aspects of queries. Then, it adopts a multi-facet retriever to build a broad pool of candidate documents covering those identified sub-aspects. However, such redundant candidates inevitably contain much irrelevant noise and are hard to handle completely by LLMs due to limited input length. As a result, sorting out the top-k most valuable documents from the candidate pool is critical to the success of the RichRAG framework.
@@
In further, we claim that a promising top-k ranking should have the following desirable features: (1) Comprehensiveness. Incentivizing LLM to generate rich and reliable responses requires external documents to comprehensively cover various query aspects. Therefore, the ability to model relationships among documents, hence maximizing the coverage of the entire reference list is crucial for the ranking module. (2) Alignment with the LLMs’ preferences. In RAG systems, the users of IR models are LLMs instead of humans. Thus, the reference order should be LLM-friendly, hence enhancing the generator to produce satisfying responses.
@@
To achieve this, we devise a generative list-wise ranker based on encoder-decoder structures. It takes as input the user query, its identified sub-aspects, and all candidates, then directly generates top-k document IDs as final ranking lists. This structure offers two key advantages: (1) Global Document Modeling. The seq-to-seq model structure equips the ranker to effectively model global interactions among candidate documents, queries, and sub-aspects, thereby capturing the overall utility of generated ranking lists in covering the query’s multi-aspects. (2) Efficiency. Following the FiD structure Izacard and Grave (2021), we parallelize the encoding of each candidate and further introduce pooling and reuse operations to the decoder module. These strategies significantly reduce the spatiotemporal load of the ranker.
@@
The optimization of our ranker consists of two stages: The first is supervised fine-tuning (SFT). To enhance the coverage of generated ranking lists on query aspects, we devise a coverage utility function based on which to build silver generation targets for training samples greedily. These silver targets allow us to SFT our ranker and ensure its basic ability. To further improve the ranking quality and align ranking lists with LLMs’ preferences, a reinforcement learning stage is introduced. We measure the LLMs’ final response quality to create reward values and adopt the DPO Rafailov et al. (2023) algorithm to optimize our ranker. In addition, we devise a unilateral significance sampling strategy (US3) to build valuable training samples for stable optimization. Experiments on two public datasets prove that RichRAG can effectively and efficiently generate more comprehensive answers for multi-faceted queries than existing methods.
@@ 2.1 Retrieval-augmented Generation
The RAG paradigm has exhibited excellent abilities in enhancing LLMs to generate accurate responses. To effectively optimize the system, some studies focus on concurrently or asynchronously training retrievers and generators Izacard et al. (2024); Borgeaud et al. (2022); Wang et al. (2023a); Arora et al. (2023); Paranjape et al. (2022); Lewis et al. (2020); Lin et al. (2023); Asai et al. (2024). However, as LLMs grow in size and inference costs rise, researchers have also explored fixing LLMs and optimizing retrievers as plug-in modules Shi et al. (2023); Yu et al. (2023) or introducing post-retrieval components, e.g., compressors and rankers Xu et al. (2023); Wang et al. (2023c); Li et al. (2024); Ke et al. (2024); Gao et al. (2024). Some studies Chan et al. (2024); Wang et al. (2024a); Khot et al. (2023); Yao et al. (2023) propose to decompose multi-hop questions into sub-tasks and solve them step-by-step to produce specific answers. However, our study mainly focuses on the query scenarios with various implicit sub-intents that require rich and comprehensive responses. It is different from the query decomposition scenario. Recent studies Rackauckas (2024) also highlight the importance of exploring users’ sub-intents, but its simplistic pipeline fails to model global document-intent interactions, leading to sub-optimal results.
@@ 2.2 Generative Ranking with LLM
Recently, the rise of LLMs allows researchers to establish various generative ranking models Nogueira et al. (2020); Zhuang et al. (2023a); Sun et al. (2023); Tamber et al. (2023); Zhuang et al. (2023b); Qin et al. (2024), including point-wise Nogueira et al. (2020); Zhuang et al. (2023a), pair-wise Zhuang et al. (2023b); Qin et al. (2024), and list-wise Tamber et al. (2023); Sun et al. (2023) models. To handle extensive document load, some generative list-wise methods Sun et al. (2023); Tamber et al. (2023) adopt a sliding-window approach to iteratively generate final ranking lists. Some studies Ke et al. (2024); Xu et al. (2024) also explore list-wise rankers in RAG systems but still focus on scenarios with specific intents and answers, neglecting the depth and breadth of user questions.
@@ 3.2 Sub-aspect Explorer
Examining various sub-aspects under a user’s query could provide explicit insights into the user’s underlying intents, thereby enabling more satisfactory results for users Santos et al. (2010); Dang and Croft (2012); Liu et al. (2020); Wang et al. (2023b). We leverage LLMs to build our sub-aspect explorer, $\mathcal{E}$, due to their extensive world knowledge and excellent capabilities in language understanding and generation. This module takes a prompt $p_{se}$, which instructs the LLM to predict the sub-aspects of the input query, and a user’s query, $q$, as input and generates a series of sub-aspects under the query, i.e., ${\hat{\mathcal{S}}}=\{\hat{s}_{1},\ldots,\hat{s}_{m}\}$, as follows:
$\{\hat{s}_{1},\ldots,\hat{s}_{m}\}=\mathcal{E}(q,p_{se}).$ (1)
@@ 3.4 Generative List-wise Ranker
Though we have collected plenty of candidates related to various sub-aspects, directly providing these massive documents to the generator is challenging due to the extensive processing burden and potential noisy information. Consequently, we devise a ranking model that targets to sort out the top-$k$ most valuable documents from the candidate pool. These ranked documents should collectively cover the query’s various sub-aspects and adhere to the preferences of the generator, hence enhancing the response performance. To equip our ranker with the ability to globally model relationships among candidates, we build it upon a generative model, T5 Raffel et al. (2020), which views all candidates, sub-aspects, and the query as input and directly generates a top-k ranking list of document IDs (docids). For each candidate, $d_{i}$, we concatenate it with the original query, $q$, its associated sub-aspects, $s(d_{i})=\{\hat{s}^{i}_{1},\ldots,\hat{s}^{i}_{n_{i}}\}$, and some special tokens to formulate an input sequence:
@@
where $r(i)$ project the rank position $i$ into the generated docid that is ranked at the $i$-th position. This pooling operation could significantly reduce the time-space burden of the decoder.
@@ 3.4.1 Supervised Fine-tuning
The silver target list, $L^{*}_{k}$, allows us to supervised fine-tune the ranker via the NTP task:
@@ 4.1 Datasets and Metrics
Metrics. To measure the matching scores of models’ responses with long-form ground truth answers, we select Rouge-1, Rouge-2, and Rouge-L as evaluation metrics. [^2] https://pypi.org/project/rouge/ Furthermore, we leverage com-rouge scores, which is introduced in Eq. 11, to assess the coverage of responses on sub-answers. We implement $\phi()$ in Eq. 11 by Rouge-1,-2, and -L to build Com-Rouge-1, -2, and -L evaluation metrics. For briefness, R1, R2, RL, CR1, CR2, and CRL are utilized to represent these metrics.
@@ 4.3 Overall Results
To fully evaluate the effectiveness of our proposed framework, RichRAG, we utilize two settings to conduct experiments, the first one we provide the predicted sub-aspects to retriever and ranker while in the second one, we provide the golden sub-aspects to unlock the RichRAG’s powers in the fullest extent possible. We present the overall results in Table 1 and have the following conclusions:
Settings Models WikiPassageQA WikiAsp
R1 R2 RL CR1 CR2 CRL R1 R2 RL CR1 CR2 CRL
Predicted Sub-aspects Close-book .1350 .0324 .0989 .2052 .0397 .1714 .0913 .0114 .0581 .1189 .0128 .0834
No Ranker .2637 .1191 .2065 .3406 .1310 .2969 .1528 .0479 .0923 .1813 .0502 .1257
RankT5 .2719 .1313 .2123 .3486 .1441 .3027 .1529 .0481 .0932 .1813 .0504 .1263
LDIST .2729 .1325 .2140 .3497 .1447 .3044 .1528 .0479 .0923 .1996 .0624 .1376
LiT5 .2735 .1291 .2118 .3501 .1413 .3033 .1723 .0594 .1037 .2013 .0620 .1399
RAG-Fusion .2583 .1146 .2029 .3355 .1261 .2913 .1628 .0547 .0976 .1912 .0572 .1324
+RankT5 .2673 .1278 .2122 .3439 .1402 .3015 .1551 .0507 .0939 .1829 .0532 .1277
+LDIST .2669 .1253 .2114 .3445 .1394 .3022 .1711 .0587 .1021 .1996 .0612 .1380
+LiT5 .2649 .1226 .2071 .3418 .1347 .2987 .1696 .0589 .1017 .1978 .0616 .1373
BGM .2637 .1191 .2065 .3373 .1537 .2983 .1485 .0472 .0920 .1764 .0497 .1242
RichRAG .2930 .1538 .2316 .3688 .1664 .3225 .1839 .0678 .1094 .2127 .0706 .1458
Golden Sub-aspects No Ranker .3400 .1667 .2650 .4164 .1809 .3600 .1956 .0655 .1277 .2259 .0688 .1671
RankT5 .3535 .1760 .2734 .4270 .1887 .3673 .1923 .0632 .1244 .2228 .0665 .1637
LDIST .3543 .1766 .2728 .4296 .1906 .3680 .2144 .0792 .1354 .2441 .0826 .1767
LiT5 .3595 .1844 .2808 .4349 .1991 .3724 .2151 .0801 .1385 .2446 .0834 .1795
RAG-Fusion .3512 .1764 .2740 .4263 .1901 .3687 .2129 .0790 .1376 .2424 .0825 .1762
+RankT5 .3576 .1825 .2774 .4333 .1972 .3702 .2035 .0718 .1289 .2332 .0751 .1693
+LDIST .3616 .1851 .2787 .4358 .1996 .3734 .2164 .0791 .1366 .2459 .0825 .1767
+LiT5 .3646 .1906 .2842 .4393 .2053 .3776 .2165 .0800 .1369 .2457 .0833 .1770
BGM .3393 .1686 .2723 .4161 .1824 .3651 .1922 .0642 .1282 .2231 .0677 .1666
RichRAG .3973 .2247 .3055 .4688 .2392 .4015 .2321 .0942 .1456 .2606 .0976 .1863
Models WikiPassageQA
R1 R2 RL CR1 CR2 CRL
RichRAG .3973 .2247 .3055 .4688 .2392 .4015
w/o SA .3262 .1505 .2610 .4046 .1648 .3516
w/o GLR .3467 .1706 .2703 .4229 .1850 .3638
w/o RL .3854 .2100 .2946 .4584 .2248 .3907
w/o US3 .3853 .2102 .2954 .4583 .2251 .3910
Models WikiAsp
R1 R2 RL CR1 CR2 CRL
RichRAG .2321 .0942 .1456 .2606 .0976 .1863
w/o SA .2096 .0678 .1479 .2058 .0541 .1557
w/o GLR .1742 .0504 .1190 .2441 .0806 .1737
w/o RL .2275 .0884 .1422 .2566 .0919 .1824
w/o US3 .2189 .0833 .1381 .2482 .0867 .1773
Table 2: Ablation results of RichRAG on two datasets.
@@
(2) Compared to list-wise ranking algorithms, our method still illustrate better performance. Though there exist several list-wise ranking algorithms in the RAG community, such as LiT5 and BGM, these algorithms do not explicitly model the interactions among candidates from the perspective of comprehensiveness of the user intent coverage. Without such explicit guidance, these algorithms may be trapped in a locally optimal solution, hence impacting the overall quality of ranking lists.
@@ 4.4 Ablation Studies
In order to evaluate the role of our key modules, we further conduct the following ablation studies with results presented in Table 2.
@@
Figure 3: Subset experiments with different sub-aspect amounts.
4.5 Impact of Sub-aspect Amount
@@ 4.6 Efficiency Analysis of Ranking Algorithms
Due to limited space, we provide further analysis studies in Appendix C,  D and  E.

### v3

@@ 1 Introduction
Though some advanced techniques Jiang et al. (2023); Asai et al. (2024); Wang et al. (2023d); Li et al. (2024) have been proposed, existing studies primarily focus on addressing specific problems that require concise and definitive answers. However, user intents are complex and multi-faceted, necessitating rich and comprehensive answers. As Figure 1 shows, when a user inquires about rapping-related information, a rich response about various aspects of rapping, such as origins, characteristics, and evolution could lead to a more satisfactory user experience than a superficial description.
@@
Our research is focused on developing effective RAG approaches to handle these more complex user needs. We propose a RAG framework, RichRAG, which is designed to offer diverse external knowledge that comprehensively covers various sub-aspects of multi-faceted queries, thereby enhancing the downstream generator (an LLM) to yield rich responses. RichRAG first employs a sub-aspect explorer to explicitly predict sub-aspects of queries. Then, it adopts a multi-facet retriever to build a broad pool of candidate documents covering those identified sub-aspects. However, such redundant candidates inevitably contain much irrelevant noise and are hard to handle completely by LLMs due to limited input length. As a result, sorting out the top-k best documents from the candidate pool is critical to the success of the RichRAG framework.
@@
In further, we claim that a promising top-k ranking should have the following desirable features: (1) Comprehensiveness. Incentivizing LLM to generate rich and reliable responses requires external documents to comprehensively cover various query aspects. Therefore, instead of predicting each document’s relevance independently, the ranking model should consider the relationship among documents to enhance the global coverage of the ranking list for query aspects. (2) Alignment with the LLMs’ preferences. In RAG systems, the users of IR models are LLMs instead of humans. Thus, the reference order should be LLM-friendly, hence enhancing the generator to produce satisfying responses.
@@
To achieve this, we devise a generative list-wise ranker based on encoder-decoder structures. It takes as input the user query, its identified sub-aspects, and all candidates, then directly generates top-k document IDs as final ranking lists. This structure offers two key advantages: (1) Global Document Modeling. The seq-to-seq model structure equips the ranker to effectively model global interactions among candidates, queries, and sub-aspects, thereby capturing the overall utility of generated ranking lists in covering the query’s multi-aspects. (2) Efficiency. Following the FiD structure Izacard and Grave (2021), we parallelize the encoding of each candidate and further introduce pooling and reuse operations to the decoder module. These strategies significantly reduce the spatiotemporal load of the ranker.
@@
The optimization of our ranker consists of two stages: The first is supervised fine-tuning (SFT). To enhance the coverage of generated ranking lists on query aspects, we devise a coverage utility function based on which to build silver generation targets (ranking lists of document IDs) for training samples greedily. These silver targets allow us to SFT our ranker and ensure its basic ability. To further improve the ranking quality and align ranking lists with LLMs’ preferences, a reinforcement learning stage is introduced. We consider both the accuracy and comprehensiveness of generated responses to create reward values, and adopt the DPO Rafailov et al. (2023) algorithm to optimize our ranker. In addition, we devise a unilateral significance sampling strategy (US3) to build valuable training samples for stable optimization. Experiments on two public datasets prove that RichRAG can effectively and efficiently generate more comprehensive answers for multi-faceted queries than existing methods.
@@ 2.1 Retrieval-augmented Generation
To ensure the effectiveness of RAG systems, previous studies mainly optimized retrievers and generators simultaneously Izacard et al. (2024); Borgeaud et al. (2022); Wang et al. (2023a); Arora et al. (2023); Paranjape et al. (2022); Lewis et al. (2020); Lin et al. (2023); Asai et al. (2024). Recent researchers also explored fixing LLMs and optimizing retrievers as plug-in modules Shi et al. (2023); Yu et al. (2023) or introducing post-retrieval components, e.g., compressors and rankers Xu et al. (2023); Wang et al. (2023d); Li et al. (2024); Ke et al. (2024); Gao et al. (2024), to reduce the training cost. Some studies Chan et al. (2024); Wang et al. (2024a); Khot et al. (2023); Yao et al. (2023); Xu et al. (2024a) propose to decompose multi-hop questions into sub-tasks and solve them step-by-step. These works typically focus on breaking down complex questions with clearly stated user intents into simpler questions. Our research, however, addresses a different scenario where user questions are broad and encompass various potential sub-aspects not explicitly stated. Answering such questions usually requires integrating diverse relevant information to these underlying sub-aspects to fully respond to the user’s potential information needs. Recent studies Rackauckas (2024) also highlight the importance of exploring users’ sub-intents, but its simplistic pipeline fails to model global document-intent interactions, leading to sub-optimal results.
@@ 2.2 Generative Ranking with LLM
Recently, the rise of LLMs allows researchers to establish various generative ranking models Nogueira et al. (2020); Zhuang et al. (2023a); Sun et al. (2023); Tamber et al. (2023); Zhuang et al. (2023b); Qin et al. (2024), including point-wise Nogueira et al. (2020); Zhuang et al. (2023a), pair-wise Zhuang et al. (2023b); Qin et al. (2024), and list-wise Tamber et al. (2023); Sun et al. (2023) models. To handle extensive document load, some generative list-wise methods Sun et al. (2023); Tamber et al. (2023) adopt a sliding-window approach to iteratively generate final ranking lists. Some studies Ke et al. (2024); Xu et al. (2024b) also explore list-wise rankers in RAG systems but still focus on scenarios with specific intents and answers, neglecting the depth and breadth of user questions.
@@ 3.2 Sub-aspect Explorer
Examining various sub-aspects under a user’s query could provide explicit insights into the user’s underlying intents, thereby enabling more satisfactory results for users Santos et al. (2010); Dang and Croft (2012); Liu et al. (2020); Wang et al. (2023b). We leverage LLMs to build our sub-aspect explorer, $\mathcal{E}$, due to their extensive world knowledge and excellent capabilities in language understanding and generation. This module takes a prompt $p_{se}$, which instructs the LLM to predict the sub-aspects of the input query, and a user’s query, $q$, as input and generates a series of sub-aspects under the query:
$\vspace{-2pt}{\hat{\mathcal{S}}}=\{\hat{s}_{1},\ldots,\hat{s}_{m}\}=\mathcal{E}(q,p_{se}).$ (1)
@@ 3.4 Generative List-wise Ranker
Though we have collected plenty of candidates related to various sub-aspects, directly providing these massive documents to the generator is challenging due to the extensive processing burden and potential noisy information. Consequently, we devise a ranking model that targets to sort out the top-$k$ most valuable documents from the candidate pool. These ranked documents should collectively cover the query’s various sub-aspects and adhere to the preferences of the generator, hence enhancing the response performance. To equip our ranker with the ability to globally model relationships among candidates, we build it upon a generative model, T5 Raffel et al. (2020), which views all candidates, sub-aspects, and the query as input and directly generates a top-k ranking list of document ID (docid) tokens. For each candidate, $d_{i}$, we concatenate it with the original query, $q$, its associated sub-aspects, $s(d_{i})=\{\hat{s}^{i}_{1},\ldots,\hat{s}^{i}_{n_{i}}\}$, and some special tokens to formulate an input sequence:
@@
where $r(i)$ project the rank position $i$ into the index of the document ranked at the $i$-th position. This pooling operation could significantly reduce the time-space burden of the decoder.
@@ 3.4.1 Supervised Fine-tuning
The silver target list, $L^{*}_{k}$, allows us to supervise fine-tune the ranker via the NTP task:
@@
$p(d^{*}_{t}|q,\mathcal{P},\hat{\mathcal{S}},d^{*}_{1:t-1})$ is the generation probability of the $t$-th target docid conditioned on the current question, the candidate pool, the sub-aspects, and the previously target documents. It is calculated by our ranking module.
@@ 3.4.2 Reinforcement Learning
Settings Models WikiPassageQA WikiAsp
F1 R2 RL BS CR2 CRL F1 R2 RL BS CR2 CRL
Predicted Sub-aspects Close-book .2132 .0324 .0989 .8417 .0397 .1714 .2037 .0114 .0581 .7967 .0128 .0834
No Ranker .3372 .1191 .2065 .8491 .1310 .2969 0.3029 .0479 .0923 .8073 .0502 .1257
RankT5 .3502 .1313 .2123 .8536 .1441 .3027 0.3079 .0481 .0932 .8063 .0504 .1263
LDIST .3487 .1325 .2140 .8523 .1447 .3044 .3377 .0479 .0923 .8114 .0624 .1376
LiT5 .3473 .1291 .2118 .8514 .1413 .3033 .3341 .0594 .1037 .8125 .0620 .1399
RAG-Fusion .3316 .1146 .2029 .8467 .1261 .2913 .3219 .0547 .0976 .8116 .0572 .1324
+RankT5 .3448 .1278 .2122 .8510 .1402 .3015 .3123 .0507 .0939 .8091 .0532 .1277
+LDIST .3400 .1253 .2114 .8494 .1394 .3022 .3327 .0587 .1021 .8122 .0612 .1380
+LiT5 .3386 .1226 .2071 .8491 .1347 .2987 .3344 .0589 .1017 .8137 .0616 .1373
BGM .3465 .1191 .2065 .8547 .1537 .2983 .2969 .0472 .0920 .8061 .0497 .1242
RichRAG .3638 .1538 .2316 .8549 .1664 .3225 .3610 .0678 .1094 .8194 .0706 .1458
Golden Sub-aspects No Ranker .3763 .1667 .2650 .8544 .1809 .3600 .3325 .0655 .1277 .8116 .0688 .1671
RankT5 .3854 .1760 .2734 .8569 .1887 .3673 .3329 .0632 .1244 .8107 .0665 .1637
LDIST .3867 .1766 .2728 .8558 .1906 .3680 .3609 .0792 .1354 .8163 .0826 .1767
LiT5 .3926 .1844 .2808 .8558 .1991 .3724 .3564 .0801 .1385 .8165 .0834 .1795
RAG-Fusion .3808 .1764 .2740 .8554 .1901 .3687 .3658 .0790 .1376 .8187 .0825 .1762
+RankT5 .3866 .1825 .2774 .8572 .1972 .3702 .3511 .0718 .1289 .8155 .0751 .1693
+LDIST .3894 .1851 .2787 .8610 .1996 .3734 .3690 .0791 .1366 .8190 .0825 .1767
+LiT5 .3915 .1906 .2842 .8567 .2053 .3776 .3676 .0800 .1369 .8190 .0833 .1770
BGM .3723 .1686 .2723 .8537 .1824 .3651 .3296 .0642 .1282 .8104 .0677 .1666
RichRAG .4174 .2247 .3055 .8637 .2392 .4015 .3951 .0942 .1456 .8256 .0976 .1863
@@ 4.1 Datasets and Metrics
Metrics. To measure the matching scores of models’ responses with long-form ground truth answers, we select F1, Rouge(-2 and -L), and BERT-Score as evaluation metrics.Furthermore, we leverage the com-rouge score, which is introduced in Eq. 11, to assess the coverage of responses on sub-answers. We implement $\phi()$ in Eq. 11 by Rouge-2, and -L to build Com-Rouge-2 and -L evaluation metrics. For briefness, F1, R2, RL, BS, CR2, and CRL are utilized to represent these metrics. Following Wang et al. (2023c), we also leverage GPT-4 to conduct a pairwise evaluation of our method with baseline models to confirm its effectiveness further. To comprehensively evaluate the effectiveness of our proposed ranking algorithm, we also provide the ranking performance comparison and analysis in Appendix C
Figure 3: Results of the GPT-4-based evaluation comparing our method with baseline models.
@@ 4.3 Overall Results
To fully evaluate the effectiveness of our proposed framework, RichRAG, we utilize two settings to conduct experiments, the first one we provide the predicted sub-aspects to retriever and ranker while in the second one, we provide the golden sub-aspects to unlock the RichRAG’s powers in the fullest extent possible. We present the overall results in Table 1 and Figure 3, and have the following conclusions:
@@
(2) Compared to list-wise ranking algorithms, our method still illustrates better performance. Though there exist several list-wise ranking algorithms in the RAG community, such as LiT5 and BGM, these algorithms do not explicitly model the interactions among candidates from the perspective of comprehensiveness of the user intent coverage. Without such explicit guidance, these algorithms may be trapped in a locally optimal solution, hence impacting the overall quality of ranking lists.
@@ 4.4 Ablation Studies
In order to evaluate the role of our key modules, we further conduct the following ablation studies with results presented in Figure 4. This figure shows the decline degrees of the ablation models compared with the complete RichRAG.
@@ 4.5 Efficiency Analysis of Ranking Algorithms
Due to limited space, we provide further analysis studies in Appendix C,  D,  E,  F and  G.